OpenStudy (anonymous):

Help me solve using elimination method? 2x - 5y = 7 3x - 10y = 13

OpenStudy (campbell_st):

label the equations 2x - 5y = 7 .......(1) 3x - 10y = 13 ...(2) to eliminate y multiply equation (1) by -2 and add to equation 2 -4x + 10y = -14 3x - 10y = 13 ---------------- -x = -1 divide both sides by -1 to find x.... now substitute your answer for x into either equation to find y hope it helps

OpenStudy (jayden):

2x-5y=73-10y=13 2x-5y+5y=73x-10y+5y 2x=73x-5y 2x-2x=73x-2x-5y 71x-5y=13 i think that is right

OpenStudy (jiteshmeghwal9):